OpenStudy (anonymous):

Algebra II question. So today my teacher is giving us something new to cover an I can't seem to get my grasp around it so can anyone help me? The worksheet asks us to solve the system of constraint and find the value of x and y that maximize the objective function. How would I go about solve a question like this? I tried graphing but I do not know where to go from there.

OpenStudy (dumbcow):

this is linear programming...this may help explain it for you http://www.purplemath.com/modules/linprog.htm

OpenStudy (anonymous):

\[Constraints \left[ x \ge 0, ~y \ge 0, ~y \le \frac{1}{2}x + 2,~ 5\ge y + x \right]\] \[5\ge y + x \Rightarrow y \le 5 -x\] My objective function is \(c = 7x - 3y\).  Is that anything like I'm supposed to do?

OpenStudy (dumbcow):

yes now shade in region defined by inequalities plug in points at corners of shaded region

OpenStudy (anonymous):

(0, 2), (5, 0), (0, 0) Do I plug in those points into the objective function? \[c=7x−3y \Rightarrow c = 7(0)-3(2) \Rightarrow c = -6\] \[c=7x−3y \Rightarrow c = 7(5)-3(0) \Rightarrow c = 35\] \[c=7x−3y \Rightarrow c = 7(0)-3(0) \Rightarrow c = 0\]

OpenStudy (dumbcow):

yes, but also the point where the lines cross 5-x = 1/5x + 2

OpenStudy (dumbcow):

of course since 7x-3y will only get smaller when y gets larger....th max occurs at point (5,0)

OpenStudy (anonymous):

I think I get it :) Thank you so much sir!

OpenStudy (dumbcow):

yw :)